# A New Action for Superstring Field Theory

### v1

@@ 1 Introduction
String field theory (SFT) gives a second-quantised formulation of string theory that facilitates the discussion of non-perturbative phenomena such as classical solutions and tachyon condensation. Its quantisation reproduces the correlation functions of 1st-quantised string theory while allowing a systematic treatment of IR divergences, renormalisation, unitarity and UV finiteness. See Sen:2024nfd for a recent review with an extensive list of references. The issue of diffeomorphism symmetry and how it is related to the symmetries of SFT together with the question of how gravity is formulated in SFT will be of particular interest here and has been addressed in Ghoshal:1991pu ; Hull:2009mi ; Mazel:2025fxj ; Mamade:2025kfk ; Mamade:2025jbs .
@@
However, the superstring field theory depends on a choice of $c=15$ superconformal field theory (SCFT) (e.g. the 10 free bosons $X^{\mu}$ and 10 free fermions $\psi^{\mu}$ for for NSR formulation of the superstring in $D=10$ Minkowski space) and whereas the equation of motion for the physical superstring field is independent of this choice of background Sen:2017szq , the action does depend on the choice of background SCFT. For an SCFT defined by a non-linear sigma model, the dependence on the SCFT will involve a dependence on a background metric, given by the target space metric of the sigma model. Secondly, it was shown in Mamade:2025jbs that the SFT gauge symmetry leads to gauge transformations with a vector parameter that act somewhat like diffeomorphisms on some fields, but under which other fields are invariant and yet others transform into each other. These transformations were referred to as ‘exotic diffeomorphisms’ in Mamade:2025jbs . However, all fields must transform under changes of coordinates and so all must transform under diffeomorphisms. As the exotic diffeomorphisms leave some fields invariant, they appear to be distinct from genuine diffeomorphisms and the question arises as to whether the SFT has proper diffeomorphism symmetry and so whether the theory is independent of the choice of coordinates. This is related to fact that, in this theory, gravity is not universal in the sense that there are fields that do not couple to the physical graviton; these are the ones that are inert under the exotic diffeomorphisms. Moreover, the SFT gauge algebra closes on field-dependent transformations and the same is expected to be true for the exotic diffeomorphisms (although some simplification is to be hoped for after field redefinitions) Mamade:2025jbs . On the other hand, general relativity is expected to be part of the low energy effective action of string theory, and the above features make it hard to see general relativity might emerge from SFT.
This superstring field theory led to Sen’s action Sen:2015nph ; Sen:2019qit for self-dual gauge fields, which also involves a doubling of the fields. This theory was inspired by the SFT construction and it has recently been conformed that it arises directly from the SFT Mamade:2025jbs . This theory gives the physical gauge field $A_{4}$ which couples to gravity and the other physical fields, together with a 2nd gauge field $\widetilde{A}_{4}$ which also has self-dual field strength but which decouples from the physical fields and in particular from $A_{4}$ and from gravity. Instead, $\widetilde{A}_{4}$ couples only to a Minkowski metric. There is a diffeomorphism-like symmetry under which the graviton transforms in the standard way but under which $\widetilde{A}_{4}$ is invariant while some fields have non-standard transformations in that they transform into other fields. These transformations are part of the ‘exotic diffeomorphisms’ of Mamade:2025jbs .
@@
In this paper, a new SFT action for the superstring will be presented which incorporates some of the features found in the bi-metric action for self-dual gauge fields. The new SFT again contains two decoupled sectors, but now both are superstrings with a full set of self-interactions. The Sen SFT is recovered when the coupling constant for the self-interactions of the shadow superstring is set to zero, so that it becomes a free theory, while the physical sector remains fully interacting. However, with interactions in both sectors, the action is independent of the choice of background and has two exotic diffeomorphism symmetries, one for each sector. The diagonal subgroup acts on all fields and is related to conventional diffeomorphism symmetry. The construction will be motivated by an analysis of the effective actions for the gravitons that arise. General relativity achieves a background-independent formulation of a graviton field $h$ in a background spacetime $\bar{g}$ by depending only on the total metric $g=\bar{g}+h$ and, in a similar spirit, the SFT presented here is background independent in the sense that it can be viewed as depending only on total string fields and not separately on a background and fluctuations.
@@ 3 Free Graviton Action
$S_{grav}=\tfrac{1}{2}\int d^{10}x\left(h^{\mu\nu}\partial^{2}h_{\mu\nu}+2(\partial^{\nu}h_{\mu\nu})^{2}\right)\,$ (17)
with the SFT gauge invariance giving rise to the linearised diffeomorphism symmetry
@@
$S_{grav}=\tfrac{1}{2}\int d^{10}x\left(h^{\mu\nu}\partial^{2}h_{\mu\nu}+2(\partial^{\nu}h_{\mu\nu})^{2}\right)-\tfrac{1}{2}\int d^{10}x\left(\hat{h}^{\mu\nu}\partial^{2}\hat{h}_{\mu\nu}+2(\partial^{\nu}\hat{h}_{\mu\nu})^{2}\right)$ (22)
where
@@
and $\hat{h}$ can be viewed as a graviton mode for the string field $\widehat{\Psi}=\Psi-{\cal G}\widetilde{\Psi}$. It was seen that the string fields $\Psi,\widehat{\Psi}$ satisfy decoupled equations of motion which, to linear order, are simply $Q\Psi=0$, $Q\widehat{\Psi}=0$, so that it was to be expected that there should be two gravitons, one from $\Psi$ and one from $\widehat{\Psi}$. The field equation for both $h_{\mu\nu}$ and $\hat{h}_{\mu\nu}$ (and also for $\tilde{h}_{\mu\nu}$) is the linearised Einstein equation. However, the action for the graviton $h$ has the correct sign for a unitary theory while that for $\hat{h}$ has the opposite sign. This action has two gauge invariances, one with parameter $\lambda_{\mu}$ for which $h$ is the gauge field and one with parameter $\hat{\lambda}_{\mu}$ for which $\hat{h}$ is the gauge field:
@@ 4.2 Gravitons in a Background Spacetime
$S_{grav}=\tfrac{1}{2}\int d^{10}x\sqrt{-\bar{g}}\left(h^{\mu\nu}\bar{\nabla}^{2}h_{\mu\nu}+2(\bar{\nabla}^{\nu}h_{\mu\nu})^{2}\right)-\tfrac{1}{2}\int d^{10}x\sqrt{-\bar{g}}\left(\hat{h}^{\mu\nu}\bar{\nabla}^{2}\hat{h}_{\mu\nu}+2(\bar{\nabla}^{\nu}\hat{h}_{\mu\nu})^{2}\right)$ (25)
where indices are raised, lowered and contracted using the background metric $\bar{g}_{\mu\nu}$, while $\bar{\nabla}$ is the Levi-Civita connection for $\bar{g}_{\mu\nu}$. The form of the first term follows from graviton correlation functions in the 1st quantised approach, while symmetry between the two sectors and the requirement that the action reduces to (17) in Minkowski space then determines the second term.
@@ 4.3 Background Independence for Einstein Gravity
The free graviton action
$S_{grav}=\tfrac{1}{2}\int d^{10}x\sqrt{-\bar{g}}\left(h^{\mu\nu}\bar{\nabla}^{2}h_{\mu\nu}+2(\bar{\nabla}^{\nu}h_{\mu\nu})^{2}\right)$ (28)
can be promoted to a background-independent action by adding interactions. The Einstein-Hilbert action
@@ 4.4 Background Independence for the Physical Sector SFT Gravitons
$S=S_{EH}(\bar{g}+\kappa h)-\tfrac{1}{2}\int d^{10}x\sqrt{-\bar{g}}\left(\hat{h}^{\mu\nu}\bar{\nabla}^{2}\hat{h}_{\mu\nu}+2(\bar{\nabla}^{\nu}\hat{h}_{\mu\nu})^{2}\right)$ (35)
@@
4.5 Background Independence for All SFT Gravitons
There is a natural way to regain background independence and hence a relation between the gauge symmetries and diffeomorphisms. The quadratic action for $\hat{h}$ can be made background independent by adding non-linear terms in $\hat{h}$ to construct the Einstein Hilbert action
@@ 5.1 Background Dependence and Interactions
to the free action to obtain the Sen action (4) obtained background independence for the field equation
@@ 5.2 The New Superstring Field Theory Action
It follows that the transformation of $\widehat{\Psi}$ is
@@
Then $\Psi$ transforms under the symmetries with parameter $\Lambda$, $\widehat{\Psi}$ transforms under the symmetries with parameter $\widehat{\Lambda}$ and $\widetilde{\Psi}$ transforms under both symmetries..
@@ 6.1 Background Independence in String Field Theory
This is to be contrasted with the notion of background independence discussed in Sen:2017szq ; Sen:1990hh ; Sen:1993mh ; Sen:1993kb ; Sen:1994kx . Any $c=15$ SCFT constitutes a consistent background for classical superstring theory and can be used to construct a superstring field theory, while any $c=26$ CFT constitutes a consistent background for the bosonic string and can be used to construct a bosonic string field theory. A longstanding question is whether different choices of (S)CFT can be regarded as giving rise to the same string field theory. Two (S)CFTs $\cal C$} and $\cal C$$^{\prime}$ will give rise to string field theories $\cal S$} and $\cal S$$^{\prime}$ respectively. The first question is whether the string background defined by $\cal C$$^{\prime}$ can be identified with a solution of the string field theory $\cal S$}. If it can, then there are two string field theories for the background defined by $\cal C$$^{\prime}$: there is $\cal S$} expanded about the solution and there is $\cal S$$^{\prime}$. The key question is whether these are the same string field theories, related by a field redefinition. For the bosonic string, this was analysed in Sen:1993mh ; Sen:1993kb for the case in which the $c=26$ CFT $\cal C$$^{\prime}$ is given by an infinitesimal marginal deformation of the CFT $\cal C$} and it was shown that the two string field theories are indeed equivalent. This then implies that any two CFTs related by a finite marginal deformation also define the same string field theory and so the string field theory can be said to be locally independent of the background used to define it.333 In comparing the state spaces of different CFTs, it was necessary to introduce a connection on the space of CFTs Sen:2017szq ; Sen:1993mh ; Sen:1993kb ; Ranganathan:1992nb ; Ranganathan:1993vj .
@@ 6.2 Background Independence for the New Superstring Field Theory Action
In this subsection we will set $\kappa=\hat{\kappa}=1$. Let
@@
defines a map ${\cal H}_{c}^{n}\to\widetilde{\cal H}_{c}$. The deformed BRST charge $Q_{\bar{\Psi}}$ and brackets $A_{1},A_{2},\dots A_{n}]_{\bar{\Psi}}$ satisfy the same algebraic conditions as $Q$ and $A_{1},A_{2},\dots A_{n}]$ and so again define a twisted $L_{\infty}$ algebra. It will also be useful to define a deformed BRST operator acting on $\widetilde{\cal H}_{c}$, with $\widetilde{Q}_{\bar{\Psi}}:\widetilde{\cal H}_{c}\to\widetilde{\cal H}_{c}$ defined by
@@ 7 The Heterotic String
The construction of Sen’s heterotic string field theory is similar to that of the type II SFT, and again uses two string fields, $\Psi$ and $\widetilde{\Psi}$, both annihilated by (2). It is interesting that using a first-quantised approach to the heterotic string using the bi-metric formulation of chiral scalars on the world-sheet also leads to a doubling of the heterotic string, with a physical heterotic string and a decoupled shadow one HullNew . The string fields are states in the Hilbert space of a $(1,0)$ SCFT with $c_{L}=15,c_{R}=26$, coupled to ghosts and bosonised superghosts. The supersymmetric left-moving sector has picture number 1-1 for the NS part of both string fields, picture number 1/2-1/2 for the R sector of the $\Psi$ field, and picture number 3/2-3/2 for the R sector of the $\widetilde{\Psi}$ field. Then
@@ 8 Discussion
In the discussion of background independence of the new SFT action, we have focussed on background solutions of the form (85) and have set $\kappa=\hat{\kappa}$. This corresponds to constructing the full theory as an expansion about a single SCFT. More generally, for a general background and for $\kappa$ and $\hat{\kappa}$ unrelated, the two sectors of the theory are expanded about different backgrounds, corresponding to different SCFTs with different BRST charges and brackets. One is used in the field equation for $\psi$ and the other for the field equation for $\hat{\psi}$. This is the analogue of the gravitational action (45) expanded about two different background metrics (54). As in that case, the resulting theory has two shift symmetries. The more general action based on two SCFTs will be discussed elsewhere.

### v2

@@ 1 Introduction
String field theory (SFT) gives a second-quantised formulation of string theory that facilitates the discussion of non-perturbative phenomena such as classical solutions and tachyon condensation. Its quantisation reproduces the correlation functions of 1st-quantised string theory while allowing a systematic treatment of IR divergences, renormalisation, unitarity and UV finiteness. See Sen:2024nfd for a recent review with an extensive list of references. Further reviews include Erbin:2021smf ; Maccaferri:2023vns ; Erler:2019loq . The issue of diffeomorphism symmetry and how it is related to the symmetries of SFT together with the question of how gravity is formulated in SFT will be of particular interest here and has been addressed in Ghoshal:1991pu ; Hull:2009mi ; Mazel:2025fxj ; Mamade:2025kfk ; Mamade:2025jbs .
@@
However, the superstring field theory depends on a choice of $c=15$ superconformal field theory (SCFT) (e.g. the 10 free bosons $X^{\mu}$ and 10 free fermions $\psi^{\mu}$ for for NSR formulation of the superstring in $D=10$ Minkowski space) and whereas the equation of motion for the physical superstring field is independent of this choice of background Sen:2017szq , the action does depend on the choice of background SCFT. For an SCFT defined by a non-linear sigma model, the dependence on the SCFT will involve a dependence on a background metric, given by the target space metric of the sigma model. Secondly, it was shown in Mamade:2025jbs that the SFT gauge symmetry leads to gauge transformations with a vector parameter that act somewhat like diffeomorphisms on some fields, but under which other fields are invariant and yet others transform into each other. These transformations were referred to as ‘exotic diffeomorphisms’ in Mamade:2025jbs . However, all fields must transform under changes of coordinates and so all must transform under diffeomorphisms. As the exotic diffeomorphisms leave some fields invariant, they appear to be distinct from genuine diffeomorphisms and the question arises as to whether the SFT has proper diffeomorphism symmetry and so whether the theory is independent of the choice of coordinates. This is related to fact that, in this theory, gravity is not universal in the sense that there are fields that do not couple to the physical graviton; these are the ones that are inert under the exotic diffeomorphisms. Moreover, the SFT gauge algebra closes on field-dependent transformations and the same is expected to be true for the exotic diffeomorphisms (although some simplification is to be hoped for after field redefinitions) Mamade:2025jbs . On the other hand, general relativity is expected to be part of the low energy effective action of string theory, and the above features make it hard to see how general relativity might emerge from this SFT.
This superstring field theory led to Sen’s action Sen:2015nph ; Sen:2019qit for self-dual gauge fields, which also involves a doubling of the fields. This theory was inspired by the SFT construction and it has recently been confirmed that it arises directly from the SFT Mamade:2025jbs . This theory gives the physical gauge field $A_{4}$ which couples to gravity and the other physical fields, together with a second gauge field $\widetilde{A}_{4}$ which also has self-dual field strength but which decouples from the physical fields and in particular from $A_{4}$ and from gravity. Instead, $\widetilde{A}_{4}$ couples only to a Minkowski metric. There is a diffeomorphism-like symmetry under which the graviton transforms in the standard way but under which $\widetilde{A}_{4}$ is invariant, while some fields have non-standard transformations in that they transform into other fields. These transformations are part of the ‘exotic diffeomorphisms’ of Mamade:2025jbs .
@@
In this paper, a new SFT action for the superstring will be presented which incorporates some of the features found in the bi-metric action for self-dual gauge fields. The new SFT again contains two decoupled sectors, but now both are superstrings with a full set of self-interactions. The Sen SFT is recovered when the coupling constant for the self-interactions of the shadow superstring is set to zero, so that it becomes a free theory, while the physical sector remains fully interacting. However, with interactions in both sectors, the action is independent of the choice of background and has two exotic diffeomorphism symmetries, one for each sector. The diagonal subgroup acts on all fields and is related to conventional diffeomorphism symmetry. The construction will be motivated by an analysis of the effective actions for the gravitons that arise. General relativity achieves a background-independent formulation of a graviton field $h$ in a background spacetime $\bar{g}$ by depending only on the total metric $g=\bar{g}+h$ and, in a similar spirit, the SFT presented here is background independent in the sense that it can be viewed as depending only on total string fields and not separately on a background and fluctuations. For the action given here, both sectors are expanded about the same background, but the full theory is independent of this choice of background. This can be generalised to an action in which each sector is expanded around a different background; this is discussed in section 8.
@@ 3 Free Graviton Action
$S_{lin}[h,\eta]=\tfrac{1}{2}\int d^{10}x\left(h^{\mu\nu}\partial^{2}h_{\mu\nu}+2(\partial^{\nu}h_{\mu\nu})^{2}-\,h\partial^{2}h+2h\,\partial_{\mu}\partial_{\nu}\,h^{\mu\nu}\right)\,$ (17)
where $h=\eta^{\mu\nu}h_{\mu\nu}$. The SFT gauge invariance gives rise to the linearised diffeomorphism symmetry
@@
$S_{grav}=S_{lin}[h,\eta]-S_{lin}[\hat{h},\eta]$ (22)
where $S_{lin}$ is defined by (17) and $S_{lin}[\hat{h},\eta]$ is given by replacing $h$ with $\hat{h}$ in (17) with
@@
so that $\hat{h}$ can be viewed as a graviton mode for the string field $\widehat{\Psi}=\Psi-{\cal G}\widetilde{\Psi}$. It was seen that the string fields $\Psi,\widehat{\Psi}$ satisfy decoupled equations of motion which, to linear order, are simply $Q\Psi=0$, $Q\widehat{\Psi}=0$, so that it was to be expected that there should be two gravitons, one from $\Psi$ and one from $\widehat{\Psi}$. The field equation for both $h_{\mu\nu}$ and $\hat{h}_{\mu\nu}$ (and also for $\tilde{h}_{\mu\nu}$) is the linearised Einstein equation. However, the action for the graviton $h$ has the correct sign for a unitary theory while that for $\hat{h}$ has the opposite sign. This action has two gauge invariances, one with parameter $\lambda_{\mu}$ for which $h$ is the gauge field and one with parameter $\hat{\lambda}_{\mu}$ for which $\hat{h}$ is the gauge field:
@@ 4.2 Gravitons in a Background Spacetime
$S_{grav}=S_{lin}[h,\bar{g}]-S_{lin}[\hat{h},\bar{g}]$ (25)
where
$S_{lin}[h,\bar{g}]=\tfrac{1}{2}\int d^{10}x\sqrt{-\bar{g}}\left(h^{\mu\nu}\bar{\nabla}^{2}h_{\mu\nu}+2(\bar{\nabla}^{\nu}h_{\mu\nu})^{2}-\,h\bar{\nabla}^{2}h+2h\,\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}\,h^{\mu\nu}\right)\,$ (26)
Here indices are raised, lowered and contracted using the background metric $\bar{g}_{\mu\nu}$, while $h=\bar{g}^{\mu\nu}h_{\mu\nu}$ and $\bar{\nabla}$ is the Levi-Civita connection for $\bar{g}_{\mu\nu}$. The form of the first term $S_{lin}[h,\bar{g}]$ follows from graviton correlation functions in the 1st quantised approach, while symmetry between the two sectors and the requirement that the action reduces to (17) in Minkowski space then determines the second term $S_{lin}[\hat{h},\bar{g}]$.
@@ 4.3 Background Independence for Einstein Gravity
The free graviton action $S_{lin}[h,\bar{g}]$ given by (26) can be promoted to a background-independent action by adding interactions. The Einstein-Hilbert action
@@ 4.4 Background Independence for the Physical Sector SFT Gravitons
$S=S_{EH}(\bar{g}+\kappa h)-S_{lin}[\hat{h},\bar{g}]$ (35)
@@
4.5 Background Independence for the SFT Gravitons
There is a natural way to regain background independence and hence a relation between the gauge symmetries and diffeomorphisms. The quadratic action $S_{lin}[\hat{h},\bar{g}]$ for $\hat{h}$ can be made background independent by adding non-linear terms in $\hat{h}$ to construct the Einstein Hilbert action
@@ 5.1 Background Dependence and Interactions
to the free action to obtain the Sen action (4) attained background independence for the field equation
@@ 5.2 The New Superstring Field Theory Action
The proof of gauge invariance uses the generalised Jacobi identities of the twisted $L_{\infty}$ algebra; the details will be given in HullPrep . It follows that the transformation of $\widehat{\Psi}$ is
@@
Then $\Psi$ transforms under the symmetries with parameter $\Lambda$, $\widehat{\Psi}$ transforms under the symmetries with parameter $\widehat{\Lambda}$ and $\widetilde{\Psi}$ transforms under both symmetries.
@@ 6.1 Background Independence in String Field Theory
This is to be contrasted with the notion of background independence discussed in Sen:2017szq ; Sen:1990hh ; Sen:1993mh ; Sen:1993kb ; Sen:1994kx . Any $c=15$ SCFT constitutes a consistent background for classical superstring theory and can be used to construct a superstring field theory, while any $c=26$ CFT constitutes a consistent background for the bosonic string and can be used to construct a bosonic string field theory. A longstanding question is whether different choices of (S)CFT can be regarded as giving rise to the same string field theory. Two (S)CFTs $\cal C$} and $\cal C$$^{\prime}$ will give rise to string field theories $\cal S$} and $\cal S$$^{\prime}$ respectively. The first question is whether the string background defined by $\cal C$$^{\prime}$ can be identified with a solution of the string field theory $\cal S$}. If it can, then there are two string field theories for the background defined by $\cal C$$^{\prime}$: there is $\cal S$} expanded about the solution and there is $\cal S$$^{\prime}$. The key question is whether these are the same string field theories, related by a field redefinition. For the bosonic string, this was analysed in Sen:1993mh ; Sen:1993kb for the case in which the $c=26$ CFT $\cal C$$^{\prime}$ is given by an infinitesimal marginal deformation of the CFT $\cal C$} and it was shown that the two string field theories are indeed equivalent. This then implies that any two CFTs related by a finite marginal deformation also define the same string field theory and so the string field theory can be said to be locally independent of the background used to define it. [^3] In comparing the state spaces of different CFTs, it was necessary to introduce a connection on the space of CFTs Sen:2017szq ; Sen:1993mh ; Sen:1993kb ; Ranganathan:1992nb ; Ranganathan:1993vj .
@@ 6.2 Background Independence for the New Superstring Field Theory Action
In this subsection we will set $\kappa=\hat{\kappa}=1$ for simplicity of presentation; the general case will be discussed in HullPrep . Let
@@
defines a map ${\cal H}_{c}^{n}\to\widetilde{\cal H}_{c}$. The deformed BRST charge $Q_{\bar{\Psi}}$ and the deformed brackets $A_{1},A_{2},\dots A_{n}]_{\bar{\Psi}}$ satisfy the same algebraic conditions as $Q$ and $A_{1},A_{2},\dots A_{n}]$ and so again define a twisted $L_{\infty}$ algebra. It will also be useful to define a deformed BRST operator acting on $\widetilde{\cal H}_{c}$, with $\widetilde{Q}_{\bar{\Psi}}:\widetilde{\cal H}_{c}\to\widetilde{\cal H}_{c}$ defined by
@@ 7 The Heterotic String
The construction of Sen’s heterotic string field theory is similar to that of the type II SFT, and again uses two string fields, $\Psi$ and $\widetilde{\Psi}$, both annihilated by (2). It is interesting that using a first-quantised approach to the heterotic string using the bi-metric formulation of chiral scalars on the world-sheet also leads to a doubling of the heterotic string, with a physical heterotic string and a decoupled shadow one Hull:2025rxy . The string fields are states in the Hilbert space of a $(1,0)$ SCFT with $c_{L}=15,c_{R}=26$, coupled to ghosts and bosonised superghosts. The supersymmetric left-moving sector has picture number 1-1 for the NS part of both string fields, picture number 1/2-1/2 for the R sector of the $\Psi$ field, and picture number 3/2-3/2 for the R sector of the $\widetilde{\Psi}$ field. Then
@@ 8 Discussion
In the new SFT action presented here, both sectors are constructed about the same background. However, it is natural to ask whether it is possible to use a different background for each sector. In the discussion of background independence of the new SFT action, we have focussed on background solutions of the form (85) and have set $\kappa=\hat{\kappa}$. This corresponds to constructing the full theory as an expansion of both sectors about a single SCFT. As has been seen, this theory has the important feature that it depends just on the two total string fields, not serperately on the background and the fluctuations. However, these two total string fields can then each be expanded about a different background, with two different coupling constants, $\kappa,\hat{\kappa}$. These two backgrounds correspond to two different SCFTs with different BRST charges and brackets. One is used in the field equation for $\psi$ and the other for the field equation for $\hat{\psi}$. This is the analogue of the gravitational action (45) expanded about two different background metrics (54). This gives a more general SFT action based on two SCFTs and will be discussed in detail in HullPrep .